OpenStudy (anonymous):

Two balls undergo inelastic collision. The y-momentum after the collision is 98 kilogram meters/second, and the x-momentum after the collision is 100 kilogram meters/second. What is the magnitude of the resultant momentum after the collision?

OpenStudy (anonymous):

you can solve this in two parts. so apply Pythagoras. \[\sqrt{x^2+y^2}\] =\[\sqrt{98^2+100^2}\] =9800 kilograms/meters second.